OpenStudy (anonymous):

A pyramid has a volume of 16 and an altitude of 5. What is the area of its base?

OpenStudy (anonymous):

1/3*b*h

OpenStudy (anonymous):

so it would be 15?

OpenStudy (anonymous):

Well, using MITL3ARN3R's formula: 1/3*b*h = 16 b*h = 48 since height is 5, we plug that in. b*5= 48 b=48/5 The area of the base would be 48/5, or 9.6 units squared. (: